OpenStudy (anonymous):

Solve t = p+3r/6 for r i tried solving it but im not sure if i got the answer right. p/6-t/6=3r ???

OpenStudy (anonymous):

t - p = 3r/6 t - p = r/2, multiply by 2: 2(t-p) = r. t = (p+3r)/6, multiply by 6 6t = p + 3r, subtract p 6t-p = 3r, divide by 3: (6t-p)/3 = r.

OpenStudy (anonymous):

thnk you